# The role of peculiar velocity uncertainties in standard siren cosmology

### v1

@@ 5 Results
As an initial illustration of the effects of fully marginalising over peculiar velocity uncertainties, in Fig. 1 we display the $H_{0}$ posteriors for Scenario 1 (neglecting linear-theory velocity covariance) and Scenario 2 (including velocity covariance) for a single realisation with a fiducial set of survey configuration parameters: $N=10$ sirens distributed to $D_{\rm max}=100$ Mpc with $5\%$ distance errors. The result indicates that the $H_{0}$ posterior is broadened by the inclusion of the peculiar velocity covariance, with the error in $H_{0}$ increasing from $1.7\%$ to $2.2\%$. Fig. 1 also demonstrates that the recovered $H_{0}$ posteriors are consistent with the input value $H_{0,{\rm fid}}=70\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$.
Figure 1: The $H_{0}$ posterior probability distribution for a bright siren analysis of $N=10$ sirens distributed to $D_{\rm max}=100$ Mpc with $5\%$ distance errors. The narrower black posterior represents the result of an analysis neglecting velocity covariance (Scenario 1 as presented in Sec. 4), whilst the broader red posterior corresponds to including the full effects of the velocity covariance (Scenario 2). The difference in the $H_{0}$ error between these two cases gets larger as the siren distance and distance error reduce.
@@
Figure 2: The recovered values of $H_{0}$ as a function of the number of sirens observed. We fix the maximum distance from the observer $D_{\rm max}=100$ Mpc and the fractional error in the distance $f_{D}=0.05$, and measure $H_{0}$ for $N=[2,5,10,25,100]$. Results from $H_{0}$ posteriors derived for a set of independent sirens including reconstruction (Scenario 1) are shown in black, results for correlated sirens including reconstruction (Scenario 2) are shown in red, and results for correlated sirens excluding reconstruction (Scenario 3) are shown in blue. The fiducial value of $H_{0}=70$ km s-1 Mpc-1 for the simulation is shown as a dotted line.
@@
Figure 3: The fractional error in measurements of $H_{0}$ as a function of the number of sirens observed. We fix the maximum distance from the observer $D_{\rm max}=100$ Mpc and the fractional error in the distance $f_{D}=0.05$, and measure $\sigma_{H_{0}}$ for $N=[2,5,10,25,100]$. Results from $H_{0}$ posteriors derived for a set of independent sirens (Scenario 1) are given by the solid black lines, results from sets of correlated sirens including reconstruction (Scenario 2) are given by the dashed red lines, and results for correlated sirens excluding reconstruction (Scenario 3) are given by the dot-dashed blue lines.
@@
Figure 4: The fractional error in measurements of $H_{0}$ as a function of the maximum distance from the observer. We fix the number of bright sirens $N=10$ and the fractional error in the distance $f_{D}=0.05$, and measure $\sigma_{H_{0}}$ for $D_{\rm max}=[50,100,150,200]$ Mpc. The curves are labelled in the same style as Fig. 3.
@@
From our exploration of the survey configuration parameter space we hence find that, whilst the introduction of correlated velocity errors has an impact in all scenarios, the relative effects are greatest for smaller survey distances and smaller distance measurement errors. Neglecting the information from velocity-field reconstructions further exacerbates these differences. The error in $H_{0}$ resulting from our simulated analyses lies in broad agreement with other bright-siren forecasts (e.g., Chen et al., 2018; Mortlock et al., 2019; Feeney et al., 2019), when consistent settings are chosen for input assumptions such as the magnitude of distance errors, and their scaling with the signal-to-noise of the input data or distance. These previous studies neglect the velocity covariance, but perform more sophisticated modelling of the gravitational-wave dataset that is beyond the scope of our analysis.
@@ Acknowledgements
This research was conducted by the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (project number CE230100016) and funded by the Australian Government.

### v2

@@ 5 Results
As an initial illustration of the effects of fully marginalising over peculiar velocity uncertainties, in Fig. 1 we display the $H_{0}$ posteriors for Scenario 1 (neglecting linear-theory velocity covariance) and Scenario 2 (including velocity covariance) for a single realisation with a fiducial set of survey configuration parameters: $N=10$ sirens distributed to $D_{\rm max}=100$ Mpc with SNR-scaled distance errors. The result indicates that the $H_{0}$ posterior is broadened by the inclusion of the peculiar velocity covariance, with the error in $H_{0}$ increasing from $1.2\%$ to $1.8\%$. Fig. 1 also demonstrates that the recovered $H_{0}$ posteriors are consistent with the input value $H_{0,{\rm fid}}=70\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$.
Figure 1: The $H_{0}$ posterior probability distribution for a bright siren analysis of $N=10$ sirens distributed to $D_{\rm max}=100$ Mpc with SNR-scaled distance errors. The narrower black posterior represents the result of an analysis neglecting velocity covariance (Scenario 1 as presented in Sec. 4), whilst the broader red posterior corresponds to including the full effects of the velocity covariance (Scenario 2). The difference in the $H_{0}$ error between these two cases gets larger as the siren distance and distance error reduce.
@@
Figure 2: The recovered values of $H_{0}$ as a function of the number of sirens observed. We fix the maximum distance from the observer $D_{\rm max}=100$ Mpc, use SNR-scaled distance errors, and measure $H_{0}$ for $N=[2,5,10,25,100]$. Results from $H_{0}$ posteriors derived for a set of independent sirens including reconstruction (Scenario 1) are shown in black, results for correlated sirens including reconstruction (Scenario 2) are shown in red, and results for correlated sirens excluding reconstruction (Scenario 3) are shown in blue. The fiducial value of $H_{0}=70$ km s-1 Mpc-1 for the simulation is shown as a dotted line.
@@
Figure 3: The fractional error in measurements of $H_{0}$ as a function of the number of sirens observed. We fix the maximum distance from the observer $D_{\rm max}=100$ Mpc, and use SNR-scaled distance errors, and measure $\sigma_{H_{0}}$ for $N=[2,5,10,25,100]$. Results from $H_{0}$ posteriors derived for a set of independent sirens (Scenario 1) are given by the solid black lines, results from sets of correlated sirens including reconstruction (Scenario 2) are given by the dashed red lines, and results for correlated sirens excluding reconstruction (Scenario 3) are given by the dot-dashed blue lines.
@@
Figure 4: The fractional error in measurements of $H_{0}$ as a function of the maximum distance from the observer. We fix the number of bright sirens $N=10$, and use SNR-scaled distance errors, and measure $\sigma_{H_{0}}$ for $D_{\rm max}=[50,100,150,200]$ Mpc. The curves are labelled in the same style as Fig. 3.
@@
From our exploration of the survey configuration parameter space we hence find that, whilst the introduction of correlated velocity errors has an impact in all scenarios, the relative effects are greatest for smaller survey distances and smaller distance measurement errors. Neglecting the information from velocity-field reconstructions further exacerbates these differences. The error in $H_{0}$ resulting from our simulated analyses lies in broad agreement with other bright-siren forecasts (e.g., 2018Natur.562..545C; 2019PhRvD.100j3523M; 2019PhRvL.122f1105F), when consistent settings are chosen for input assumptions such as the magnitude of distance errors, and their scaling with the signal-to-noise of the input data or distance. These previous studies neglect the velocity covariance, but perform more sophisticated modelling of the gravitational-wave dataset that is beyond the scope of our analysis.
@@ Acknowledgements
We thank an anonymous referee for providing useful comments on the original draft of this paper. This research was conducted by the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (project number CE230100016) and funded by the Australian Government.